E-time + D-time 20150402 UT

Observers: JHill, DMiller, OKuhn, BRothberg (Tucson)
Telescope Operator: GBechetti (LBTO)
AO Support: DMiller (Tucson)
Software Support: (Tucson)

Summary

Plan

  • PEPSIPFU and GCS tests twilight until 03:30
  • Barry science with LUCI1 3:30 until 4:30
  • Doug's crosstalk measurements 4:30 until 6:30
  • Barry and Olga science 7 until 10
  • WRS and LBC tests 10 until twilight

Preparations

Installed GCS patch GCS.20150401 for Issue 5521.

Details

PEPSIPFU on both sides

02:15 ACQUIRE preset to Beta Tau - some funniness about X polynomial required restarting PCS. SX IE=-22 CA=-31 DX IE=-5 CA=-208

02:32 ACTIVE preset - did not find SX star -- manual ACQUIRE -- SX IE=+3 SX CA=-39 DX IE=-6 DX CA=-208

02:51 ACTIVE preset DX star moved by ~10 arcsec. Made Issue 5525 about the peculiar pointing.

02:52 perturbed SX guiding with ptincrement 0 IE 2 and it responded perfectly. Issue 5521 is closed.

Sky is clear with seeing varying from 0.6 to 1.1 arcsec.

Ilya et al. are testing spilled light guiding on the PFU fiber with this V=7.2 star (it seems to be working).

Science with LUCI1

03:34 handover to Barry

See LBTO notes in the Partner Observing wiki

Crosstalk check

05:30 Collect z4,11,22 on right
  • R cross_1 Some Z22 and Z4 on the Z11 request

06:00 Collect z4,11,22 on left
  • having some troubles collimating to less than 300 nm RMS. start with LIMIT=450.
  • L cross_1

06:09 Change Z11/22 from 0.2 to 0.3 and Z11/4 from 5.0 to 5.5 (same as LEFT)
  • R cross_2

06:20 R collimate_multiple, 10, /send_zern, /send_cow, EXP=15.0
  • R equ_1 through wfsc 55
  • for wfsc 49 change rdg_recon.list to have all 0.0 for rotation
  • 06:25 send 5 more collimations

06:29 L collimate_multiple, 10, /send_zern, /send_cow, EXP=15.0
  • L equ_1
  • Diverge!

06:39 R collimate_multiple, 10, /send_zern, /send_cow, EXP=15.0
  • but first offset 30, 0
  • R equ_2

06:40 on LBTO account L Z4, 11, 22 to see if new idl routines cause problem (last night on dlmiller account was old, now new routines)
  • L equ_1

06:42 offset, 60, 0
  • R equ_3

06:55 offset 90, 0
  • R egu_4 Crap. Guiding is jumping. telescope moved through 88+ elevation. my bad!

07:09 preset to AO1001 in NE

07:13 no offset, collimate_multiple 10, RIGHT, equ_5
  • send_equ 85-94

07:19 equ_5
  • send_sext 95-104

07:22 Left has been collimating on-axis
  • on-axis GCS through 266
  • offset, 45, 0
  • GCS continue to collimate through 277 (?)
  • offset, 90, 0
  • GCS continue to collimate through 299
  • offset, 120, 0

07:24 offset, 45, 0 equ_5
  • send_equ 105-114

07:32 equ_5
  • send_sext 115-124

07:38 offset, 90, 0, equ_5
  • send_equ 125-134

07:44 equ_5
  • send_sext 135-144

08:30 R Changed Z11/4 from 5.0 to 5.5
  • R cross_3 Z4 is better. Set Z11/4 to 5.8. Did not collect a data set. Closed for wind

-- %USERSIG{JohnHill - 2015-04-02}%
Topic revision: r3 - 02 Apr 2015, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback